40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  Cryo Lab eLog  Not logged in ELOG logo
Message ID: 2686     Entry time: Thu Mar 18 16:44:33 2021
Author: shruti 
Type: DailyProgress 
Category: PSOMA 
Subject: cavity alignment, RF oscillator power and frequency 

Proceeded to align the cavity as outlined in elog 2685.

 

I also measured the output of OCXO using the Moku; all channels have 10dB attenuation and 50 Ohm impedance.

Attachment 1: 32.7 MHz; Channel 1 (red) is 'TO EOM', Channel 2 (blue) is 'TO LO'.

Attachment 2: 33.59 MHz; Channel 1 (red) is 'TO EOM', Channel 2 (blue) is 'TO LO'.

Attachment 3: Channel 1 (red) is 'TO LO' at 32.57 MHz, Channel 2 (blue) is 'TO LO' at 33.59 MHz.

The data files for all the spectra can be found here.

 

 

Attachment 1: MokuSpectrumAnalyzerData_20210318_162955_Screenshot.png  925 kB  Uploaded Thu Mar 18 17:50:35 2021  | Hide | Hide all
MokuSpectrumAnalyzerData_20210318_162955_Screenshot.png
Attachment 2: MokuSpectrumAnalyzerData_20210318_162552_Screenshot.png  916 kB  Uploaded Thu Mar 18 17:50:51 2021  | Hide | Hide all
MokuSpectrumAnalyzerData_20210318_162552_Screenshot.png
Attachment 3: MokuSpectrumAnalyzerData_20210318_162358_Screenshot.png  881 kB  Uploaded Thu Mar 18 17:51:13 2021  | Hide | Hide all
MokuSpectrumAnalyzerData_20210318_162358_Screenshot.png
ELOG V3.1.3-